Calculate the osmotic pressure of a solution containing 24.6 g of glycerin (C3H8O3) in 266.5 mL of solution at 310 K .
1
Expert's answer
2020-03-10T09:07:32-0400

Molar mass of glycerine = 92 g

moles of glycerine = 24.6/92 = 0.27

Molarity"=\dfrac{moles}{volume(in \ litres)}=\dfrac{0.27}{0.2665}"="1"

Osmotic pressure "(\pi)=CRT" where C is the molarity R is the gas constant and T is the temperature in kelvin

"\pi= 1*0.082*310=25.5 atm"